OpenStudy (anonymous):

Find the unit normal vector to the surface at the given point. z=(x^2+y^2)^1/2 (3,4,5)

OpenStudy (anonymous):

\[z=\sqrt{x^2+y^2}, at the point (3,4,5)\]

OpenStudy (anonymous):

the same of finding gradient at the given point

OpenStudy (anonymous):

find partial derivatives, and take their value at your point. That's the vector normal to the surface

OpenStudy (anonymous):

\[Gradient( z)= (x/\sqrt{x ^{2}+y ^{2}},y/\sqrt{x ^{2}+y ^{2}})\]

OpenStudy (anonymous):

Calculating the partials, I got : \[0=z-\sqrt{x^2+y^2}\] \[\Delta F(x,y,z) = x/\sqrt{x^2+y^2}i+y/\sqrt{x^2+y^2}j-1k\]

OpenStudy (anonymous):

^ yeah thats what I got.

OpenStudy (anonymous):

so now just put your values in

OpenStudy (anonymous):

Then when I evaluated the gradient I get: \[\Delta F(3,4,5) = <3\sqrt{5}/5,4\sqrt{5}/5>\]

OpenStudy (anonymous):

i guess, i didn't make the calculation

OpenStudy (anonymous):

ΔF(3,4,5)=<35 √ /5,45 √ /5,1>

OpenStudy (anonymous):

how are you getting 35 and 45?

OpenStudy (amistre64):

gradient isnt \Delta, its \nabla\[\nabla F ...\]